# Inference for multiple treatment effects using confounder importance learning

### v4

@@ 2. Modelling Framework
We remark that much of the Bayesian treatment literature does not consider treatment inclusion indicators, rather they are forced into the model. While that strategy is easily accommodated in our framework by setting $\delta_{t}=1$ for all $t$, we consider that one often wishes to assess whether the treatment effects exist in the first place, and otherwise shrink their estimates towards zero. Accordingly with this goal, we adopt the so-called product moment (pMOM) non-local prior of Johnson and Rossell (2012). Briefly, non-local priors improve the rates at which one discards the truly zero parameters, see Johnson and Rossell (2012); Wu (2016); Shin et al. (2018); Rossell (2021). Under the pMOM prior, one has $\alpha_{t}=0$ almost surely if $\delta_{t}=0$, and
@@
with the analogous setting for every $\beta_{j}$. Figure 8 illustrates its density. Above $v_{t}$ is the sample variance of treatment $t$, to ease notation we assume that treatments and covariates have unit variance and take $v_{t}=1$. The pMOM prior involves a prior dispersion parameter $\tau>0$, that by default we set to $\tau=1/3$ following Rossell et al. (2021), which leads to a minimally informative prior akin to the unit information prior underlying the Bayesian information criterion. As for the dispersion parameter, where unknown, we also place a minimally informative $\phi\sim\text{IGam}(0.01,0.01)$ prior.
@@
The main modelling novelty in this article is the choice of covariate prior inclusion probabilities $\pi_{j}({\bm{\theta}})=P(\beta_{j}\neq 0\mid{\bm{\theta}})$, where ${\bm{\theta}}=(\theta_{0},\theta_{1},\dots,\theta_{T})$ is a key prior hyper-parameter relating $\pi_{j}({\bm{\theta}})$ to (positive) measures of association ${\bf f}_{j}=(f_{j,1},\ldots,f_{j,T})$ between covariate $j$ and the $T$ treatments. Specifically,
(4) $\pi_{j}({\bm{\theta}})=\rho+(1-2\rho)\left(1+\exp\left\{-\theta_{0}-\sum_{t=1}^{T}\theta_{t}f_{j,t}\right\}\right)^{-1}$
where $\rho\in(0,1/2)$. Model (4) is essentially that used in logistic regression where the scalar $\rho$ ensures that prior inclusion probabilities $\pi_{j}({\bm{\theta}})\in[\rho,1-\rho]$ are bounded away from (0,1), so that one does not include/exclude covariates a priori. We propose the default $\rho=1/(J^{2}+1)$, which sets a lower-bound $P(\beta_{j}\neq 0\mid{\bm{\theta}})\geq 1/(J^{2}+1)$ that is of the same order as $J$ grows as the complexity priors in Castillo and van der Vaart (2012). These bounds on the prior inclusion probabilities ensure that one may consistently recover the true model as $n\rightarrow\infty$, assuming suitable conditions, see Rossell (2021) for a detailed discussion.
Akin to DL, BAC and related methods, the idea is that if covariate $j$ is highly associated to treatment $t$ then $f_{j,t}$ will be large, and if $\theta_{t}>0$ then one favors the inclusion of such covariates. In contrast, if $\theta_{t}=0$ then said inclusion is not encouraged, and if $\theta_{t}<0$ it is discouraged. Figure 2 illustrates $\pi_{j}({\bm{\theta}})$ for three different values of $\theta_{1}$. Setting ${\bm{\theta}}$ is critical for the performance of our inferential paradigm, and in Section 3 we introduce data-driven criteria and algorithms for its choice. Intuitively, in high-confounding scenarios where covariates associated to treatment $t$ are also associated to the outcome, one expects to learn $\theta_{t}>0$. In contrast, in low-confounding scenarios where most covariates associated to treatment $t$ are instruments, one expects to learn $\theta_{t}<0$.
@@
where $\left({\mathbf{X}}^{\intercal}{\mathbf{X}}\right)^{+}$ is the Moore-Penrose pseudo-inverse, and ${\mathbf{X}}$ the $n\times J$ design matrix. For $J<n$ this is the familiar OLS estimator, but (6) is also well-defined when $J>n$, and it has been recently investigated in terms of the so-called benign overfitting property in Bartlett et al. (2020). Wang and Leng (2016) showed that when $J>n$, (6) ranks the coefficients consistently under theoretical conditions slightly broader than the LASSO. Therefore, one expects that all parents of treatment $t$ have larger values of $f_{j,t}$. Similarly, by the screening property of the LASSO one expects that all parents of treatment $t$ have $f_{j,t}=|w_{j,t}|>0$. This is appealing in our context, since $\pi_{j}({\bm{\theta}})$ are mainly driven by the relative magnitudes of $f_{j,t}$, the prior inclusion probabilities are allowed to favor or discourage the parents of treatments (depending on whether $\theta_{t}>0$ or $\theta_{t}<0$). Recall that DL, BAC and most of the literature is based on encouraging the inclusion of such covariates, which would be harmful in low-confounding scenarios.
We remark that in the ACPME framework of Wilson et al. (2018) one also pre-computes features relating treatments to covariates. Relative to BAC, pre-computing features offers a massive computational simplification, since no model search over treatment equations is needed (which would multiply the model space size by $2^{JT}$). Relative to ACPME, we measure treatment-covariate association by regressing the treatments on the controls, rather than regressing the controls on the treatments. We opted for the former option because one needs to run $T$ rather than $J$ regressions, and that as discussed this allows consistently identifying parents of the treatments as $n$ grows. Further, as discussed next, we propose learning our hyper-parameters ${\bm{\theta}}$ using likelihood-based methods, and propose scalable computational algorithms to do so. In Section 4 we show comparisons between our CIL, BAC, ACPME, and other methods.
@@
Figure 2. Prior inclusion probability (4) as a function of $f_{j,1}$, a feature measuring correlation between covariate $j$ and treatment $t=1$, for $\theta_{0}=-1$, $\rho=(J^{2}+1)^{-1}$, and $J=99$ covariates. Top and bottom dotted lines show the upper and lower bounds, $1-\rho$ and $\rho$, respectively. The dotted line in the middle corresponds to $\theta_{1}=0$.
@@ 3. Computational Methodology
3.1. Bayesian model averaging
@@
Outside particular cases such as Gaussian regression under Gaussian priors, (7) does not have a closed-form expression. To estimate (7) under our pMOM prior we adopt the approximate Laplace approximations of Rossell et al. (2021), see Section G.1 for an overview. We then obtain point estimates using BMA,
@@ 3.2. Confounder importance learning via marginal likelihood
The main challenge in obtaining ${\bm{\theta}}^{\text{EB}}$ is that evaluating $p({\mathbf{y}}\mid{\bm{\theta}})$ requires a costly sum over $({\bm{\delta}},{\bm{\gamma}})$. Fortunately, it is possible to obtain a simpler expression for the gradient of the log-marginal likelihood, given in Proposition 3.1. The proof (see Section A) leverages the fact that ${\mathbf{y}}$ is conditionally independent from ${\bm{\theta}}$ given $({\bm{\gamma}},{\bm{\delta}})$, that the prior $p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})$ factorizes, and the specific form of $\pi_{j}({\bm{\theta}})$ in (4).
@@
(9) $\nabla_{{\bm{\theta}}}\log p({\mathbf{y}}\mid{\bm{\theta}})=(1-2\rho)\sum_{j=1}^{J}\mathbf{f}_{j}\left[P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right],$
@@
Proposition 3.1 allows using gradient-based optimization to approximate ${\bm{\theta}}^{\text{EP}}$. Notice that we only need to sum over $J$ terms, as opposed to $2^{J+T}$ for evaluating the marginal likelihood, and the gradient only depends on the data via the marginal inclusion probabilities $P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$, which can typically be estimated more accurately than joint model probabilities. However, one must still compute $P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$ for every considered ${\bm{\theta}}$, which is intensive when the optimization requires more than a few iterations, since typically an MCMC algorithm will be used to estimate these probabilities. We next describe an Expectation Propagation variational approximation which, as illustrated in Figure S4, typically provides a good approximation to the global mode.
@@ 3.3. Confounder importance learning by Expectation-Propagation
The use of Expectation Propagation (EP) (Minka, 2001a, b) is common in machine learning, including in variable selection (Seeger et al., 2007; Hernández-Lobato et al., 2013; Xu et al., 2014). We propose a computationally efficient approximation to the marginal likelihood optimizer that can be used as is, or serve as initialization for a the gradient-based optimization discussed in Section 3.2. We comment on our experience with both schemes later in this section.
To simplify notation, we first define the posterior inclusion probabilities for a specific value of ${\bm{\theta}}$:
@@
the right-hand side following from the fact that, when ${\bm{\theta}}=\bm{0}$, the prior on the model space $p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}}=0)=1/2^{J+T}$ is the uniform distribution. We consider a mean-field approximation to $p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})$,
@@
The parameters that minimize the Kullback-Leibler divergence between $p_{0}$ and $q$,
(11) $(\mathbf{r}^{\textsc{ep}},\mathbf{s}^{\textsc{ep}})=\operatorname*{arg\,max}_{\mathbf{r},\mathbf{s}}\sum_{{\bm{\gamma}},{\bm{\delta}}}p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})\log\left({p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})\over q({\bm{\delta}},{\bm{\gamma}})}\right)\,.$
are given by
@@
The way we propose to use this approximation for a computationally inexpensive estimator of ${\bm{\theta}}$ is as follows. First, note that
@@
Our strategy is to replace $p_{0}$ by $q$, using the optimized variational parameters in Proposition 3.2, which reduces the above sum from one over $2^{J+T}$ terms to one over $J$ terms. Specifically, it is possible to show that (see Section E.1) this leads to
@@
Our strategy is to pre-compute $r_{j}^{\textsc{ep}}=P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}}=\mathbf{0})$, prior to conducting the optimization in (14). This leads to an optimization over ${\bm{\theta}}\in\mathbb{R}^{T+1}$ where, in contrast to the marginal likelihood estimate ${\bm{\theta}}^{\text{EB}}$, the objective function can be cheaply evaluated. Since the function in (14) is not concave, we conduct an initial grid search and subsequently use a quasi-Newton algorithm, see Algorithm 1 in the supplement. Although this was not an issue in our examples, when the number of treatments $T$ is large the mentioned grid search becomes too costly. Possible alternatives are either using Bayesian optimization methods or simply initializing at ${\bm{\theta}}=0$ (uniform model prior) and using the gradient of the objective function in (14), which we give in Section E.2.

### v6

@@ 2.1. Model
We remark that much of the Bayesian treatment effects literature does not consider treatment inclusion indicators, rather they are forced into the model. While that strategy is easily accommodated in our framework by setting $\delta_{t}=1$ for all $t$, we consider that one often wishes to assess whether the treatment effects exist in the first place, and otherwise shrink their estimates towards zero. Accordingly with this goal, we adopt the so-called product moment (pMOM) non-local prior of Johnson and Rossell (2012). Briefly, non-local priors improve the rates at which one discards the truly zero parameters, see Johnson and Rossell (2012); Wu (2016); Shin et al. (2018); Rossell (2021). Under the pMOM prior, one has $\alpha_{t}=0$ almost surely if $\delta_{t}=0$, and
@@
with the analogous setting for every $\beta_{j}$. Figure S3 illustrates its density. Above $v_{t}$ is the sample variance of treatment $t$, to ease notation we assume that treatments and covariates have unit variance and take $v_{t}=1$. The pMOM prior involves a prior dispersion parameter $\tau>0$, that by default we set to $\tau=1/3$ following Rossell et al. (2021), which leads to a minimally informative prior akin to the unit information prior underlying the Bayesian information criterion. As for the dispersion parameter, where unknown, we also place a minimally informative $\phi\sim\text{IGam}(0.01,0.01)$ prior.
@@
The main modelling novelty in this article is the choice of covariate prior inclusion probabilities $\pi_{j}({\bm{\theta}})=P(\beta_{j}\neq 0\mid{\bm{\theta}})$, where ${\bm{\theta}}=(\theta_{0},\theta_{1},\dots,\theta_{T})$ is a key prior hyper-parameter relating $\pi_{j}({\bm{\theta}})$ to (positive) measures of association ${\bf f}_{j}=(f_{j,1},\ldots,f_{j,T})$ between covariate $j$ and the $T$ treatments. Specifically, prior probabilities are given by a logistic regression equation with success probability
(4) $\tilde{\pi}_{j}({\bm{\theta}})=\left(1+\exp\left\{-\theta_{0}-\sum_{t=1}^{T}\theta_{t}f_{j,t}\right\}\right)^{-1}$
truncated to lie in a pre-specified interval $[\underline{\rho},\bar{\rho}]$. That is,
$\pi_{j}({\bm{\theta}})=\begin{cases}\underline{\rho}\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\leq\bar{\rho}\\ \tilde{\pi}_{j}({\bm{\theta}})\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})\\ \bar{\rho}\mbox{, if }\tilde{\pi}_{j}({\bm{\theta}})\geq\bar{\rho}\end{cases}.$
The truncation to $[\underline{\rho},\bar{\rho}]$ ensures that one does not include/exclude covariates a priori, and is required for the asymptotic properties discussed in Section 4. We propose default $\underline{\rho}=1/J$ and $\bar{\rho}=0.95$. The former allows enforcing sparsity, while ensuring that the prior expected model size is non-decreasing in $J$. The latter avoids assigning overly strong evidence a priori that a covariate is needed, which helps prevent covariate over-selection. See Section 4 for further discussion on $(\underline{\rho},\bar{\rho})$.
Akin to DL, BAC and related methods, the idea is that if covariate $j$ is highly associated to treatment $t$ then $f_{j,t}$ will be large, and if $\theta_{t}>0$ then one favors the inclusion of such a covariate. In contrast, if $\theta_{t}=0$ then said inclusion is not encouraged, and if $\theta_{t}<0$ it is discouraged. Figure 2 illustrates $\pi_{j}({\bm{\theta}})$ for three different values of $\theta_{1}$. Setting ${\bm{\theta}}$ is critical for the performance of our inferential paradigm, and in Section 3 we introduce data-driven criteria and algorithms for its choice. Intuitively, in high-confounding scenarios where covariates associated to treatment $t$ are also associated to the outcome, one expects to learn $\theta_{t}>0$. In contrast, in low-confounding scenarios where most covariates associated to treatment $t$ are instruments, one expects to learn $\theta_{t}<0$.
@@
where $\left({\mathbf{X}}^{\intercal}{\mathbf{X}}\right)^{+}$ is the Moore-Penrose pseudo-inverse, and ${\mathbf{X}}$ the $n\times J$ design matrix. For $J<n$ this is the familiar OLS estimator, but (6) is also well-defined when $J>n$, and it has been recently investigated in terms of the so-called benign overfitting property in Bartlett et al. (2020). Wang and Leng (2016) showed that when $J>n$, (6) ranks the coefficients consistently under theoretical conditions slightly broader than the LASSO. Therefore, one expects that all parents of treatment $t$ have larger values of $f_{j,t}$. Similarly, by the screening property of the LASSO one expects that all parents of treatment $t$ have $f_{j,t}=|w_{j,t}|>0$. This is appealing in our context, since $\pi_{j}({\bm{\theta}})$ are mainly driven by the relative magnitudes of $f_{j,t}$, the prior inclusion probabilities are allowed to favor or discourage the parents of treatments (depending on whether $\theta_{t}>0$ or $\theta_{t}<0$).
We remark that the ACPME framework of Wilson et al. (2018) also pre-computes features relating treatments to covariates. The main difference is that we use the outcome data to estimate ${\bm{\theta}}$, whereas in ACPME it is a fixed hyper-parameter. By fixing ${\bm{\theta}}$, ACPME cannot adapt the prior behavior depending on whether there is low or high confounding (or neither), which is critical to prevent variance inflation. This occurs by design, if one does not use the outcome data, one cannot assess whether controls are confounders or instruments (i.e. associated or not with the outcome). In contrast CIL seeks to estimate $\theta_{t}>0$ when there is high confounding between treatment $t$ and covariates, as measured by our novel confounding coefficient (Section 2.2). Critically, CIL can also set $\theta_{t}=0$ or even $\theta_{t}<0$ when there is no confounding to help exclude instruments, a feature that is not provided by ACPME (nor other competing methods, to our knowledge). The other main difference is that our features $f_{j,t}$ are obtained by regressing the treatments on the covariates, whereas in ACPME features are obtained by regressing the covariates on the treatments. Large $f_{j,t}$ suggests that covariate $j$ is a parent of treatment $t$, which is critical to interpret the CIL solution as setting prior probabilities that reflect the true value of the confounding coefficient (Section 4). Such an interpretation is not possible for ACPME. In Section 5 we show comparisons with BAC, ACPME, and other methods.
@@
Figure 2. Prior inclusion probability (4) as a function of $f_{j,1}$, a feature measuring correlation between covariate $j$ and treatment $t=1$, for $\theta_{0}=-1$, and $J=99$ covariates. The grey lines show the lower and upper bounds, $\underline{\rho}=1/J$ and $\bar{\rho}=0.95$, respectively.
2.2. Confounding coefficient
For each treatment $t=1,\ldots,T$, we define a confounding coefficient $\kappa_{t}^{*}$ quantifying the extent to which the treatment is truly confounded with the $J$ covariates in (1). We also define a sample-based counterpart $\kappa_{t}$. Intuitively, large positive $\kappa_{t}^{*}$ indicates high confounding, i.e. that covariates truly associated with treatment $t$ are mostly confounders ($\beta_{j}\neq 0$ in (1)), whereas large negative $\kappa_{t}^{*}$ indicates no confounding, i.e. that said covariates are mostly instruments ($\beta_{j}=0$). $\kappa_{t}^{*}=0$ indicates neutral confounding, i.e. being associated to the treatment is unrelated to being a confounder or instrument.
To provide the definition, we introduce two elements. First, let $\beta_{j}^{*}$ be the Kullback-Leibler optimal parameter in (1) quantifying the effect of covariate $j$ on the outcome (if (1.1) is correctly specified, then $\beta_{j}^{*}$ is the true covariate effect). Let ${\bm{\gamma}}^{*}=(\gamma_{1}^{*},\ldots,\gamma_{d}^{*})$, where $\gamma_{j}=\mbox{I}(\beta_{j}^{*}\neq 0)$, so that $|{\bm{\gamma}}^{*}|_{0}/J$ is the proportion of truly active covariates. Second, let $f_{j,t}^{*}$ be a measure of the (unknown) true effect of covariate $j$ on treatment $t$. For example, consider a generalized linear model where the mean of treatment $t$ is truly driven by $\sum_{j=1}^{J}w_{j,t}^{*}x_{i,j}$ and let $f_{j,t}^{*}=|w_{j,t}^{*}|$ (under model misspecification, $w_{j,t}^{*}$ can be defined as the Kullback-Leibler optimal parameter). Another interesting example is to define $f_{j,t}^{*}=\mbox{I}(w_{j,t}^{*}\neq 0)$, an indicator for covariate $j$ truly having an effect on treatment $t$. Without loss of generality, we assume that the vector $(f_{1,t}^{*},\ldots,f_{j,t}^{*})$ has zero mean and unit variance.
Definition 2.1.
The confounding coefficient $\kappa_{t}^{*}$ is
(7) $\kappa_{t}^{*}=\frac{1}{J}\sum_{j=1}^{J}f_{j,t}^{*}\left(\gamma_{j}^{*}-\frac{|{\bm{\gamma}}^{*}|_{0}}{J}\right)$
Note that $\kappa_{t}^{*}/\sqrt{V({\bm{\gamma}}^{*})}$ is the correlation between covariate effect indicators ${\bm{\gamma}}^{*}$ and treatment-covariate association features $(f_{1,t}^{*},\ldots,f_{j,t}^{*})$. Hence, when $\kappa_{t}^{*}>0$, large $f_{j,t}^{*}$ (covariate $j$ is associated with treatment $t$) indicates that it is likely that $\gamma_{j}^{*}=1$ (covariate $j$ is a confounder). In contrast, when $\kappa_{t}^{*}<0$ then it is more likely that $\gamma_{j}^{*}=0$ (covariate $j$ is an instrument). Finally, we let $\kappa_{t}=\frac{1}{J}\sum_{j=1}^{J}f_{j,t}(\gamma_{j}-|{\bm{\gamma}}|_{0}/J)$ be the sample confounding coefficient, i.e. replacing $f_{j,t}^{*}$ by their estimates $f_{j,t}$ and acknowledging that ${\bm{\gamma}}^{*}$ is unknown.
@@ 3. Computational Methodology
3.1. Bayesian model averaging (BMA)
@@
Outside particular cases such as Gaussian regression under Gaussian priors, (8) does not have a closed-form expression. To estimate (8) under our pMOM prior we adopt the approximate Laplace approximations of Rossell et al. (2021), see Section S6.1 for an overview. We then obtain point estimates using BMA,
@@ 3.2. Confounder importance learning via marginal likelihood
The main challenge in obtaining ${\bm{\theta}}^{\text{EB}}$ is that evaluating $p({\mathbf{y}}\mid{\bm{\theta}})$ requires a costly sum over $({\bm{\delta}},{\bm{\gamma}})$. Fortunately, it is possible to obtain a simpler expression for the gradient of the log-marginal likelihood, given in Proposition 3.1. The proof (see Section S2) leverages the fact that ${\mathbf{y}}$ is conditionally independent from ${\bm{\theta}}$ given $({\bm{\gamma}},{\bm{\delta}})$, that the prior $p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}})$ factorizes, and the specific form of $\pi_{j}({\bm{\theta}})$ in (4).
@@
(10) $\nabla_{{\bm{\theta}}}\log p({\mathbf{y}}\mid{\bm{\theta}})=\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right].$
@@
Proposition 3.1 allows using gradient-based optimization to approximate ${\bm{\theta}}^{\text{EP}}$. Notice that we only need to sum over at most $J$ terms, as opposed to $2^{J+T}$ for evaluating the marginal likelihood. Also, the gradient only depends on the data via the marginal inclusion probabilities $P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$, which can typically be estimated more accurately than joint model probabilities. However, one must still compute $P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$ for every considered ${\bm{\theta}}$, which is intensive when the optimization requires more than a few iterations, since typically an MCMC algorithm will be used to estimate these probabilities. Section 3.3 describes an expectation-propagation variational approximation that in most of our experiments provided a good approximation to the global mode.
The empirical Bayes solution given by Proposition 3.1 has a natural interpretation. For simplicity we discuss the case where $\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})$ for all $j=1,\ldots,J$. When some $\pi_{j}({\bm{\theta}})\not\in(\underline{\rho},\bar{\rho})$ a similar interpretation holds, basically one excludes covariates with $\pi_{j}({\bm{\theta}})$ equal to $\underline{\rho}$ or $\bar{\rho}$. Setting the first entry of the gradient equal to zero gives
(11) $\sum_{j=1}^{J}\pi_{j}({\bm{\theta}})=\sum_{j=1}^{J}P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$
That is, prior inclusion probabilities are set such that the prior expected model size is equal to the posterior expected model size. This seems appealing, as the latter converges to the number of truly active covariates under suitable conditions (see Section 4).
Simple algebra shows that setting the other entries of the gradient to zero gives that $E(\kappa_{t}\mid{\bm{\theta}})=E(\kappa_{t}\mid y,{\bm{\theta}})$ for $t=1,\ldots,T$, where $E(\kappa_{t}\mid{\bm{\theta}})=J^{-1}\sum_{j=1}^{J}f_{j,t}[\pi_{j}({\bm{\theta}})-\bar{\pi}]$ is the prior expectation of the sample confounding coefficient (Definition (7)), and
(12) $E(\kappa_{t}\mid y,{\bm{\theta}})=\frac{1}{J}\sum_{j=1}^{J}f_{j,t}[P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-E(|{\bm{\gamma}}|_{0}\mid{\mathbf{y}},{\bm{\theta}})]$
its posterior expectation, where $E(|{\bm{\gamma}}|_{0}\mid{\mathbf{y}},{\bm{\theta}})=J^{-1}\sum_{j=1}^{J}P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})$. Setting the prior mean of $\kappa_{t}$ to $E(\kappa_{t}\mid y,{\bm{\theta}})$ is again appealing, since the latter converges to the true confounding coefficient $\kappa_{t}^{*}$ (Section 4). Hence, CIL seeks to set prior probabilities that reflect the true amount of confounding, in particular learning from data whether one is in a high, neutral, or no confounding scenario.
@@ 3.3. Confounder importance learning by expectation-propagation
The use of expectation-propagation (EP) (Minka, 2001a, b) is common in machine learning, including in variable selection (Seeger et al., 2007; Hernández-Lobato et al., 2013). We propose a computationally efficient approximation to the marginal likelihood optimizer that can be used as is, or serve as initialization for the gradient-based optimization in Section 3.2. For simplicity, we denote the posterior inclusion probabilities for the specific value ${\bm{\theta}}=\bm{0}$:
@@
The right-hand side arises because, when ${\bm{\theta}}=\bm{0}$, the model space prior $p({\bm{\delta}},{\bm{\gamma}}\mid{\bm{\theta}}=0)=1/2^{J+T}$ is uniform. We consider a mean-field approximation to $p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})$,
@@
The minimizer of the Kullback-Leibler divergence between $p_{0}$ and $q$,
$(\mathbf{r}^{\textsc{ep}},\mathbf{s}^{\textsc{ep}})=\operatorname*{arg\,max}_{\mathbf{r},\mathbf{s}}\sum_{{\bm{\gamma}},{\bm{\delta}}}p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})\log\left({p_{0}({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}})\over q({\bm{\delta}},{\bm{\gamma}})}\right)\,.$
is given by
@@
This approximation provides a computationally inexpensive estimator, denoted ${\bm{\theta}}^{\textsc{ep}}$. First, note that
@@
Our strategy is to replace $p_{0}$ by $q$, using the variational parameters in Proposition 3.2. Section Section S4.1 shows that the above sum over $2^{J+T}$ terms is reduced to one over $J$ terms, specifically
@@
The gradient of the objective function in (16) is
$\sum_{j:\pi_{j}({\bm{\theta}})\in(\underline{\rho},\bar{\rho})}\mathbf{f}_{j}\left[P^{EP}(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})-\pi_{j}({\bm{\theta}})\right].$
where $P^{EP}(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}})=\pi_{j}({\bm{\theta}})r_{j}^{EP}/[\pi_{j}({\bm{\theta}})r_{j}^{EP}+(1-\pi_{j}({\bm{\theta}}))(1-r_{j}^{EP})]$. This expression is analogous to (10), and allows to similarly interpret ${\bm{\theta}}^{\textsc{ep}}$ in terms of confounding coefficients (see discussion after (10)).
@@
Our strategy is to pre-compute $r_{j}^{\textsc{ep}}=P(\gamma_{j}=1\mid{\mathbf{y}},{\bm{\theta}}=\mathbf{0})$, prior to conducting the optimization in (16). This leads to an optimization over ${\bm{\theta}}\in\mathbb{R}^{T+1}$ where, in contrast to the marginal likelihood estimate ${\bm{\theta}}^{\text{EB}}$, the objective function can be cheaply evaluated. Since the function in (16) is not concave, we conduct an initial grid search and subsequently use a quasi-Newton algorithm, see Algorithm 1 in the supplement. Although this was not an issue in our examples, when the number of treatments $T$ is large the mentioned grid search becomes too costly. Possible alternatives are either using Bayesian optimization methods or simply initializing at ${\bm{\theta}}=0$ (uniform model prior) and using the gradient of the objective function in (16).